OpenStudy (anonymous):

3(q-2) + 2 = 5q - 7 - 2q NEED HELP NOW!!

OpenStudy (anonymous):

;)

OpenStudy (anonymous):

best response please ;)

OpenStudy (anonymous):

thank you Lethal for ACTUALLY being of help

OpenStudy (anonymous):

first algebra hw huh. here you are going to use the distributive property 3(q-2) + 2 will then change to 3q -6 + 2 you can add -6 and 2 3q - 4 will be your left side now lets work on the right 5q-7-2q you have like terms 3q - 7 will be your right side 3q -4 = 3q -7 + 4 +4 3q = 3q -3 -3q -3q 0q = -3 ? invalid equation, no solution

OpenStudy (anonymous):

so it is False right

OpenStudy (anonymous):

yep because \[-4 \neq -7\]

OpenStudy (anonymous):

i know it was FALSE thanks ;)

OpenStudy (anonymous):

so what is a ( wingspan )

OpenStudy (anonymous):

okay thank you to both. help with 5z = -3(z+7) ???? want to double check it